OpenStudy (anonymous):

Forty-five is 20% of what number?

OpenStudy (anonymous):

20 is 1\5th of the numner you want. Multiply 45 by 5 which is 225

OpenStudy (anonymous):

2.25 or 225 because i got 2.25

OpenStudy (anonymous):

225 is correct if u wanna check take 225*20% to get 45

OpenStudy (anonymous):

20% I mean and 225. Think about it it is obciously not going to be smaller then the original number

OpenStudy (phi):

you might want to think like this: 20% of a number can be thought of as 20% times a number the other idea to remember is 20% is "short-hand" for 20/100 (which in this case can be simplified to 1/5) so here is what the question is asking 45 is (i.e. equals) 1/5 times x (x for unknown number)

OpenStudy (anonymous):

ok thanks

OpenStudy (phi):

of course you still have to figure out the number \[ 45 = \frac{1}{5} x \] if you learn this trick: multiply both sides by 5: \[ 5\cdot 45 = 5 \cdot \frac{1}{5} x \] that simplifies to \[ 5 \cdot 45 = x\]